the local amateur football club spent $675 on tickets to a professional football game. if the club bought three fewer fifteen dollar tckets than four-fifths the number of twelve dollar tickets, how many tickets of each type did the club buy?
Answer by solve_for_x(190)

Let x represent the number of $12 tickets purchased.

Then, the number of $15 tickets is "three fewer ... than four-fifths the number" of $12 tickets.

The number of $15 tickets can then be represented as:

(4/5)x - 3

The amount collected from the sale of $12 tickets is 12 * x = 12x.

The amount collected from the sale of $15 tickets is 15 * [(4/5)x - 3] = 12x - 45.

Together, these amounts must add up to the total ticket sales of $675, which
gives the following equation:

12x + (12x - 45) = 675

Collecting like terms on the left side gives:

24x - 45 = 675

Adding 45 to both sides then gives:

24x = 720

Dividing both sides by 24 then leaves:

x = 30

This is the number of $12 tickets. You can then substitute x = 30 into "(4/5)x - 3"
to get the number of $15 tickets.